OpenStudy (anonymous):

A boat travels 20 miles up the river in same amount of time it takes to travel 28 miles down the same river. If the current is 2 miles per hour, what is speed of the boat in still water?

OpenStudy (mrnood):

speed = distance/time speed upstream = Vstill - 2 speed downstream = Vstill +2 for upstream Vstill-2 = 20/t for downstream Vstill+2 = 28/t since t is the same for both journey you can solve for Vstill

OpenStudy (anonymous):

D=RT?

OpenStudy (mrnood):

Whatever the abbreviations you use speed = distance/time, distance = speed * time, time = distance/speed (do you mean D=RT "Distance = Rate*Time"?)

OpenStudy (anonymous):

yes, I have to do the little table and then solve and stuff, but I forgot where to put the #s

OpenStudy (anonymous):

x=12! Yaaa y, I did it! lol

OpenStudy (mrnood):

Well done - that's correct!

OpenStudy (anonymous):

Can you help me with the next one?